rawsavy:

Which description matches the graph of the inequality y ≥ –x – 3? a shaded region above a dashed boundary line a shaded region below a solid boundary line a shaded region above a solid boundary line a shaded region below a dashed boundary line

jhonyy9:

do you can graphing this inequality ?

jhonyy9:

what is the difference between a dashed boundary line graph and a solid boundary line graph ? - do you know it ?

rawsavy:

i dont know

surjithayer:

consider y=-x-3 when y=0,x=-3 when when x=0,y=-3 so line passes through (-3,0) and (0,-3) again consider \[y \ge -x-3\] put x=0,y=0 in above inequality \[0\ge -3\] which is true . Hence (0,0) lies in the graph. and (0,0) is above the line y=-x-3 shaded region above a solid line . C